OpenStudy (anonymous):

If 15=-2(2t-1),what does t=?

OpenStudy (anonymous):

t=-13/4

jimthompson5910 (jim_thompson5910):

\[\large 15=-2(2t-1)\] \[\large 15=-4t+2\] \[\large 15-2=-4t\] \[\large 13=-4t\] \[\large -\frac{13}{4}=t\] \[\large t=-\frac{13}{4}\]